# Revisiting Robust RAG: Do We Still Need Complex Robust Training in the Era of Powerful LLMs?

### v1

@@
Abstract
Retrieval-augmented generation (RAG) systems often suffer from performance degradation when encountering noisy or irrelevant documents, driving researchers to develop sophisticated training strategies to enhance their robustness against such retrieval noise. However, as large language models (LLMs) continue to advance, the necessity of these complex training methods is increasingly questioned. In this paper, we systematically investigate whether complex robust training strategies remain necessary as model capacity grows. Through comprehensive experiments spanning multiple model architectures and parameter scales, we evaluate various document selection methods and adversarial training techniques across diverse datasets. Our extensive experiments consistently demonstrate that as models become more powerful, the performance gains brought by complex robust training methods drop off dramatically. We delve into the rationale and find that more powerful models inherently exhibit superior confidence calibration, better generalization across datasets (even when trained with randomly selected documents), and optimal attention mechanisms learned with simpler strategies. Our findings suggest that RAG systems can benefit from simpler architectures and training strategies as models become more powerful, enabling more scalable applications with minimal complexity.
Revisiting Robust RAG: Do We Still Need Complex Robust Training
in the Era of Powerful LLMs?
Hanxing Ding1,2thanks:   Equal contributions. Shuchang Tao11footnotemark: 1  Liang Pang1thanks:   Corresponding author Zihao Wei1,2 Liwei Chen3Kun XuHuawei Shen1,2 Xueqi Cheng1,2 1Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences  3 Kuaishou Technology {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
"Entities should not be multiplied unnecessarily."
— Occam’s razor
1 Introduction
Modern LLMs excel in various NLP tasks, such as text generation, knowledge reasoning, and question answering Ye et al. (2024); Tao et al. (2024); Zhao et al. (2023). Their contextual relevance and coherence make them valuable tools across domains. Despite their broad applications, LLMs struggle with external knowledge integration, particularly in RAG systems where poor retrieval results can lead to inaccurate or misleading responses Xu et al. (2024); Tan et al. (2024); Ding et al. (2024). This issue exposes a key limitation in the robustness of LLMs with imperfect retrieval results.
@@ 1 Introduction
Figure 1: Comparison of EM scores of different document selection strategies on the TriviaQA dataset. As the model capabilities increase, the performance gap ($\Delta$) between the best and worst strategies among the four training methods decreases from 14.65% to 4.36%.
To address this challenge, researchers have proposed various robust training methods to enhance RAG’s resilience against noisy contexts. Approaches mainly focus on selecting high-quality documents or incorporating additional adversarial loss regularization terms during training Yoran et al. (2024); Fang et al. (2024); Krueger et al. (2021); Zhu et al. (2024a); Jin et al. (2024); Zeng et al. (2024). For instance, RetRobust Yoran et al. (2024) suggests using a strategic mixture of relevant and irrelevant documents to train LLMs to withstand noisy contexts. RAAT Fang et al. (2024) and ATM Zhu et al. (2024a) introduce adversarial regularization terms to encourage consistent model performance across different document contexts, making the model less sensitive to retrieval noise.
Given the increasing capabilities of LLMs, a fundamental question emerges: do sophisticated robust training strategies remain necessary as model capacity grows? We first conduct a simple preliminary experiment to investigate this inquiry. As shown in Figure 1, we compare the EM scores of four different training strategies on the TriviaQA dataset. Notably, the performance gap ($\Delta$) between the best and worst strategies shrinks dramatically from 14.65% to 4.36% as the model capacity increases. This significant diminishing gap suggests that powerful LLMs might naturally possess enhanced robustness, potentially challenging the necessity of complex robust training techniques.
In this paper, we present a systematic investigation into the necessity of sophisticated robust training strategies as model capabilities advance. Through extensive experiments across various language models with different capabilities levels, we uncover a counterintuitive phenomenon: while models with limited capabilities benefit from high-quality document selection and complex adversarial losses, these benefits diminish substantially as model capability increases. Notably, for highly capable models, training with randomly selected or even seemingly irrelevant documents can achieve comparable or superior performance to sophisticated training strategies.
We further conduct an in-depth study on this counter-intuitive phenomenon through a systematic analysis. Through inherent confidence calibration analysis, we discover that newer powerful models naturally exhibit superior discrimination between correct and incorrect responses. Moving to external performance, we find these models demonstrate remarkable cross-dataset generalization with basic random document training. Our visualization of attention patterns further reveals the underlying mechanism: powerful models inherently learn effective attention allocation even with simple training. These findings, from internal capabilities to external performance to mechanism analysis, indicate that powerful models may naturally possess fundamental desired capabilities, potentially reducing the need for complex training strategies.
We provide insights for future RAG development, suggesting: simplified architecture design principles for powerful models, opportunities for scalable open-domain applications with minimal supervision, and theoretical perspectives on how training requirements evolve with model scaling.
@@
  • Finding: We reveal that as model capabilities increase, the benefits of sophisticated training strategies diminish substantially.
  • Rationale: Through systematic analysis, we trace this finding to powerful models’ inherent capabilities, demonstrating that they naturally possess strong confidence calibration, generalization, and effective attention mechanisms, regardless of training sophistication.
  • Insights: We provide practical guidelines for future RAG development, advocating for architectural simplification with powerful models and highlighting new theoretical perspectives on model scaling laws.
2 Background
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) enhances language model performance by incorporating external knowledge retrieval into the generation process. Given a query $q$, a retriever $\mathcal{R}$ selects a set of $k$ relevant document sets $\mathcal{D}=\{d_{1},d_{2},\dots,d_{k}\}$ from a large corpus, providing additional factual grounding for the model. These retrieved documents serve as context to improve the model’s reasoning and factual accuracy. The language model $G$ then generates an answer $a$ based on the augmented input, following the conditional probability formulation:
$G(a\mid q,\mathcal{D})=\prod_{t=1}^{T}P(a_{t}\mid a_{<t},q,\mathcal{D};\theta),$ (1)
where $a=(a_{1},a_{2},\dots,a_{T})$ represents the generated sequence, and $\theta$ denotes the model parameters. The retrieval-augmented setup enables the model to dynamically condition its responses on external information, improving its ability to generate factually consistent answers. However, RAG models remain susceptible to retrieval noise, such as irrelevant or misleading documents, which can degrade performance and lead to hallucinations.
2.2 Robust Training for RAG
Document Selection Strategy
A common strategy for robust training in RAG models is to select high-quality documents from either the golden set or top-ranked passages for fine-tuning. Traditionally, supervised fine-tuning (SFT) relies on high-quality golden data, where each query is paired with a carefully curated document containing the correct answer. This ensures that the model learns from reliable sources, leading to improved factual accuracy. Naturally, leveraging such golden documents in RAG training is a straightforward and effective approach.
However, in real-world scenarios, the golden document may not always be retrieved due to limitations in the retriever. In such cases, training only on golden passages lead to unrealistic expectations during inference. A more practical approach is to incorporate top-1 retrieved documents, which better reflect what the model will encounter at test time. Even if these documents are imperfect, training on them enhances the model’s adaptability. Further, drawing from a mixture of top-ranked, lower-ranked, or randomly sampled passages, as suggested by RetRobust Yoran et al. (2024), exposes the model to diverse retrieval conditions, improving its robustness to noisy or suboptimal results. This strategy ensures that RAG models generalize better across different retrieval settings, making them more reliable in real-world applications.
Adversarial Loss Design
To enhance the robustness of RAG models against retrieval variations, robust training for RAG introduces a general training framework that combines standard SFT with a robustness-enforcing regularization term. During training, the model learns across multiple retrieval environments, which may include golden documents, top-k retrieved results, or adversarially perturbed documents. The overall training objective consists of two components: the first part is the standard training loss across different retrieval environments, and the second part is a regularization term designed to reduce model performance sensitivity to retrieval variations. Different forms of regularization strategies are detailed in Appendix B.
3 Experimental Setups
In this section, we introduce the dataset and evaluation metrics we used to assess the robustness performance of various robust training strategies.
3.1 Datasets and Evaluation Metrics
For our experiment, we evaluate on four widely-used question answering datasets: (1) single-hop QA, including NaturalQuestions (NQ) Kwiatkowski et al. (2019) and WebQuestions Berant et al. (2013); and (2) multi-hop QA, including TriviaQA Joshi et al. (2017) and HotpotQA Yang et al. (2018). All experimental results are evaluated on their development splits using the Exact Match (EM) and F1 metrics. Detailed statistics of these datasets are listed in Appendix C.1 Table 2.
3.2 RAG Pipeline
We implement a standard two-stage RAG framework with retrieval and generation phases. For retrieval, we use Contriever Izacard et al. (2022), a BERT-based dense retriever trained with unsupervised contrastive learning. Our knowledge corpus is Wikipedia-2018, preprocessed into 100-word non-overlapping passages. We encode these passages using Contriever and index them with FAISS [^1] https://github.com/facebookresearch/faiss for efficient retrieval. During training and development, the top-20 relevant documents are retrieved for each query. For question-answering, the top-5 retrieved documents are concatenated with the query as input for answer generation. The generator prompt is detailed in Appendix C.3 Table 3.
3.3 Robust RAG Training Setups
For our experiments, we evaluate two popular robust training settings:
3.3.1 Document Selection Strategies
We explore several document selection training strategies across three categories.
For basic QA capability evaluation, we use the Base Model in a few-shot setting, where the model performs inference without RAG-specific training.
For scenarios with relevant documents, we explore four strategies:
  • RALM: Fine-tunes the model by prepending golden retrieval text. Queries without golden documents in the top-20 are excluded, leading to reduced training set size.
  • RetRobust: Following Yoran et al. (2024), this method enhances robustness by randomly selecting top-ranked, low-ranked, or random passages during training.
@@ 3.3.1 Document Selection Strategies
  • Golden Document: Selects the most relevant document containing the correct answer. If none exist in the top-20, it defaults to the top-1 document for consistency.
Model RAG Scenario HotpotQA NQ WebQuestions TriviaQA AVERAGE
@@
Llama-2-7b-chat-hf Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
@@
$\Delta$ (Worst $\rightarrow$ Best) 16.19% 12.89% 21.87% 13.19% 48.53% 28.43% 13.55% 17.66% 15.95% 13.20%
Llama-3-8B-Instruct Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
RALM 27.64 38.10 35.19 46.10 47.84 56.98 54.75 63.27 41.36 51.11
RetRobust 36.06 48.99 43.28 55.04 52.88 62.10 59.06 67.77 47.82 58.48
Top-1 Doc 36.72 49.30 44.38 56.20 54.68 62.26 60.80 68.31 49.15 59.02
Golden Doc 35.52 48.31 41.35 53.13 48.92 58.41 58.26 66.99 46.01 56.71
[…]
Irrelevant Doc 35.31 47.92 42.45 54.41 46.76 57.67 58.97 66.57 45.87 56.64
[…]
IRM 35.19 48.08 41.13 53.14 53.96 61.64 57.15 69.13 46.86 58.00
$\Delta$ (Worst $\rightarrow$ Best) 14.04% 12.53% 7.90% 5.76% 16.94% 8.46% 11.74% 10.70% 10.92% 8.42%
@@
Table 1: Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA). The row $\Delta$ (Worst $\rightarrow$ Best) indicates the performance gain achieved by the best method compared to the worst strategy among these training strategies, representing the benefit of sophisticated robust training methods.
To assess robustness against irrelevant information, we introduce two adversarial scenarios:
  • Random Document: Randomly selects a document from retrieved results, simulating unpredictable retrieval quality.
  • Irrelevant Document: Chooses a passage from another query’s retrieval results, ensuring no relevance to the current query.
3.3.2 Adversarial Loss Design
We assess two popular adversarial loss strategies:
  • RAAT. The regularization term in RAAT reduces the performance gap between the best and worst retrieval cases. By penalizing excessive performance disparity, RAAT ensures that the model remains stable even under challenging retrieval conditions, leading to improved robustness and generalization.
  • IRM. The regularization in IRM minimizes the variance in performance across different retrieval environments. By enforcing consistency, IRM mitigates sensitivity to distribution shifts, ensuring that the model performs reliably across diverse retrieval scenarios.
4 Does Sophisticated Robust Training Still Matter in Powerful Models?

### v2

# On the Diminishing Returns of Complex Robust RAG Training in the Era of Powerful LLMs

@@
Abstract.
Retrieval-augmented generation (RAG) systems traditionally employ sophisticated training strategies to enhance robustness against retrieval noise. In this work, we investigate a critical question: does the benefit of these complex robust training methods diminish as language models become more powerful? Through systematic evaluation across multiple model scales and question-answering datasets, our analysis reveals a consistent trend: the marginal robustness benefit of sophisticated training strategies decreases substantially as model capacity increases. While smaller models show significant performance improvements from complex document selection and adversarial objectives, more capable models achieve comparable or even superior performance with simpler training approaches. Further investigation demonstrates that stronger models naturally exhibit better confidence calibration, cross-dataset generalization capability, and more effective attention patterns, even under simple training regimes. These findings suggest that as foundation models evolve, the engineering effort invested in complex robust training may yield diminishing returns, indicating that simplified RAG pipelines could suffice for powerful models while maintaining competitive performance.
Large Language Models, Retrieval-augmented Generation, Robust Training
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region; December 7–10, 2025; Xi’an, Chinabooktitle: Proceedings of the 2025 Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (SIGIR-AP 2025), December 7–10, 2025, Xi’an, Chinadoi: 10.1145/3767695.3769518isbn: 979-8-4007-2218-9/2025/12ccs: Information systems Question answering
1. Introduction
Modern LLMs excel in various NLP tasks including text generation, knowledge reasoning, and question answering (Ye et al., 2024; Tao et al., 2024; Zhao et al., 2023), yet they struggle with external knowledge integration, particularly in RAG systems where poor retrieval can lead to inaccurate responses (Xu et al., 2024; Tan et al., 2024; Ding et al., 2024). To address this challenge, researchers have proposed robust training methods focusing on high-quality document selection or adversarial loss regularization (Yoran et al., 2024; Fang et al., 2024; Krueger et al., 2021; Zhu et al., 2024b; Jin et al., 2024; Zeng et al., 2024). For instance, RetRobust (Yoran et al., 2024) uses a mixture of relevant and irrelevant documents during training, while RAAT (Fang et al., 2024) and ATM (Zhu et al., 2024b) incorporate adversarial regularization terms to maintain consistent performance across varied document contexts.
@@ 1. Introduction
Figure 1. Performance comparison of different robust training strategies on the TriviaQA dataset. As model capabilities increase, the marginal robustness benefit ($\Delta$), i.e., performance gain between the best and worst strategies among the four training methods decreases from 14.65% to 4.36%.
With the rapid advancement of LLM capabilities, a critical research question arises: how does the effectiveness of sophisticated robust training strategies change as model capacity grows?
To investigate this question empirically, we conduct a preliminary analysis comparing four distinct training strategies on the TriviaQA dataset, as illustrated in Figure 1. Our results reveal a striking pattern: the performance gain $\Delta$ between the best and worst strategies decreases from 14.65% to 4.36% as model capacity scales up. This pronounced reduction in $\Delta$ suggests a law of diminishing returns, wherein the marginal benefit of complex training methods becomes increasingly negligible as model capacity expands.
In this paper, building upon this observation, we conduct a systematic investigation to validate and characterize the diminishing returns of robust RAG training as model capacity increases. We propose an evaluation framework that quantifies the marginal robustness benefit ($\Delta$)—the performance gain introduced by sophisticated document selection and adversarial training methods. Our comprehensive experiments confirm and strengthen the initial finding: as model capability increases, the marginal robustness benefit of complex training diminishes sharply. While smaller models significantly benefit from curated documents and adversarial losses, stronger models achieve comparable or even superior performance with simpler setups. Strikingly, training high-capacity models with randomly selected documents often matches or exceeds the performance of advanced robust training methods, challenging fundamental assumptions about robust training strategies.
To uncover the mechanisms behind this phenomenon, we perform a comprehensive analysis of model behavior. Our findings show that stronger models possess inherent capabilities that reduce reliance on complex training. They exhibit superior confidence calibration, accurately distinguishing correct from incorrect outputs without additional supervision. Even when trained with random documents, these models generalize well across datasets. Attention visualizations further reveal that high-capacity models naturally learn effective attention patterns under simple training, explaining the diminishing gains from sophisticated methods.
These findings reveal a shift in how robust RAG systems should be designed: as model capacity increases, simpler training and retrieval strategies become not only sufficient but preferable. This shift opens up new opportunities for scalable RAG applications, and prompts a rethinking of training priorities in the context of model scaling.
@@
  • Finding: We demonstrate that the marginal benifit from sophisticated robust training diminish significantly as model capacity increases.
  • Rationale: Through detailed analysis, we attribute this finding to intrinsic properties of stronger models-including better confidence calibration, cross-task generalization, and effective attention patterns—even under simple robust training.
2. Preliminaries
2.1. Related Work
2.1.1. Retrieval-Augmented Generation
Retrieval-Augmented Generation has been developed to enhance the reasoning and generation capabilities of LLMs by integrating external information. However, dense retrievers are not always perfect and can sometimes recall irrelevant or erroneous documents, which can degrade performance (Jiang et al., 2023; Ding et al., 2024). To mitigate this issue, adaptive retrieval techniques and robust RAG training methods have been proposed. Adaptive retrieval techniques retrieve external information only when LLMs encounter unknowns, reducing the risk of integrating misleading or incorrect information (Jiang et al., 2023; Mallen et al., 2023; Asai et al., 2023). Additionally, robust training focuses on enhancing LLMs’ response to various retrieval noises, aiming to efficiently obtain adversarial examples that improve model robustness while reducing training overhead (Fang et al., 2024; Zhu et al., 2024b; Xiang et al., 2024).
2.1.2. Robust Training for RAG Systems
The performance of LLMs in RAG systems can be compromised by irrelevant context (Cuconasu et al., 2024; Wu et al., 2024). Recently, researchers have focused on this issue and proposed effective adversarial training methods to enhance the robustness of LLMs against noisy documents. RetRobust (Yoran et al., 2024) suggests using a mix of relevant and irrelevant documents to train LLMs to withstand noisy contexts. RAAT (Fang et al., 2024) and ATM (Zhu et al., 2024b) employ adversarial training to dynamically adjust the model’s training process, aiming to maintain robustness against noise and generate accurate answers. RobustRAG (Xiang et al., 2024) proposes to enhances RAG model robustness against retrieval corruption attacks through an isolate-then-aggregate strategy to achieve certifiable robustness. Additionally, another work focus on identifying key documents from noisy contexts to facilitate effective question answering through knowledge rewriting, refinement, or filtering (Zhu et al., 2024a; Qiao et al., 2024; Jin et al., 2024; Zeng et al., 2024). However, these robust training approaches are primarily applied to small or weak LMs with fewer than 7 billion parameters. Thus, there’s an urgent need to explore whether complex robust training is still necessary to improve the robustness and generalization of bigger or stronger models when dealing with noisy contexts.
2.2. Robustness Training for RAG
2.2.1. Retrieval-augmented Adaptive Adversarial Training
To improve the robustness of retrieval-augmented language models (RALMs) against retrieval noise, RAAT (Fang et al., 2024) incorporates an adversarial loss into the standard supervised fine-tuning (SFT). The model processes one golden context and three adversarial samples at each iteration, optimizing for the most challenging perturbation. The adversarial objective follows a min-max strategy:
(1) $\mathcal{L}_{\text{max}}=\max_{da\in DA}\mathcal{L}(\theta,da(q),a),$
(2) $\mathcal{L}_{\text{ada}}=\mathcal{L}_{\text{max}}+w_{\text{reg}}\cdot|\mathcal{L}_{\text{max}}-\mathcal{L}_{\text{min}}|_{2}^{2},$
2.2.2. Multiagent Iterative Tuning Optimization
ATM (Zhu et al., 2024b) steers the generator to have a robust perspective of useful documents for question answering with the help of an auxiliary attacker agent. The generator is trained to maximize answer correctness while minimizing its sensitivity to adversarial perturbations. It receives the user query along with a document list, which may contain both relevant and fabricated information introduced by the attacker. Its objective is to generate accurate responses as long as sufficient truthful information is present while reducing the impact of misleading or fabricated content. To achieve this, the generator learns to identify and leverage relevant documents while ignoring noisy ones, regardless of whether they originate from the original retrieval $\mathcal{D}$ or adversarial perturbations $\mathcal{D}^{\prime}$. This can be formalized as maximizing the objective:
(3) $G(a\mid q,\mathcal{D}^{\prime})-\mathrm{dist}\left[G(a\mid q,\mathcal{D}),G(a\mid q,\mathcal{D}^{\prime})\right],$
2.2.3. Invariant Risk Minimization
Invariant Risk Minimization (IRM) aims to learn representations that remain stable across different environments, improving generalization under distribution shifts. To enhance the robustness of retrieval-augmented generation (RAG) across varying retrieval conditions, the V-REx objective (Krueger et al., 2021) can also be adapted to enforce risk invariance across different retrieval environments. Given retrieval environments $\mathcal{E}=\{1,\dots,m\}$, where each $e\in\mathcal{E}$ corresponds to a specific retrieval scenario (e.g., golden documents, top-$k$ retrieved, noisy retrieval), we define the empirical risk $\mathcal{R}_{e}(\theta)$ of the model parameterized by $\theta$. The training objective is formulated as:
(4) $\mathcal{R}_{\textrm{V-REx-RAG}}(\theta)=$ $\beta\;\mathrm{Var}(\{\mathcal{R}_{1}(\theta),.,\mathcal{R}_{m}(\theta)\})$
$+\sum^{m}_{e=1}\mathcal{R}_{e}(\theta),$
3. Evaluating Marginal Benefit of Robust RAG Methods
In this section, we establish a systematic evaluation framework to address a fundamental question in RAG: whether sophisticated robust training methods continue to yield substantial performance improvements as model capabilities increase. We first formalize our evaluation metric to quantify the marginal benefit provided by different robust RAG strategies, followed by a detailed categorization of current state-of-the-art robust RAG methods. We also provide an experimental setup that ensures fair comparison across model scales and dataset types.
3.1. Metric: Marginal Robustness Benefit
To measure the relative utility of sophisticated robust training methods, we introduce the Marginal Robustness Benefit, denoted as $\Delta$, which captures the relative performance gain between the best and worst robust RAG strategies:
(5) $\Delta=\frac{S_{\text{best}}-S_{\text{worst}}}{S_{\text{worst}}}$
where $S_{\text{best}}$ and $S_{\text{worst}}$ denote the performance scores (EM or F1) of the most and least effective training strategies, respectively.
By computing $\Delta$ across models of varying capacities, we quantify how the relative effectiveness of robust training evolves with model scaling. This metric helps assess whether the complexity introduced by advanced training strategies remains necessary as foundation models grow more capable.
3.2. Taxonomy of Robust RAG Methods
We categorize current robust RAG training approaches into two dimensions:
3.2.1. Document Selection Strategy
We first analyze robust RAG training methods based on document selection strategies. These approaches use different criteria to select documents for supervised fine-tuning, thereby improving model adaptability to retrieval variations:
  • Golden Document: Selects the most relevant document containing the correct answer. If none exist in the top-20 documents, use the top-1 document for consistency.
@@ 3.2.1. Document Selection Strategy
  • RALM: Fine-tunes the model by prepending golden retrieval text. Queries without golden documents in the top-20 docs are excluded.
  • RetRobust: Following Yoran et al. (2024), this method enhances robustness by randomly selecting top-ranked, low-ranked, or random passages for training.
To assess robustness against irrelevant information, we also include two adversarial scenarios:
  • Random Document: Randomly selects a document from retrieved results.
  • Irrelevant Document: Uses a passage from another query’s results, ensuring complete irrelevance.
3.2.2. Adversarial Loss Design
Next, we analyze robust RAG methods based on adversarial loss designs. These approaches enhance model resilience against retrieval noise by combining standard SFT with robustness-enforcing regularization. Models train across multiple retrieval environments (golden documents, top-k retrievals, or adversarially perturbed documents) using a dual objective: standard training loss plus regularization to reduce sensitivity to retrieval variations.
Dataset Type # Train # Dev
NQ single-hop 79,168 8,757
WebQuestions single-hop 2,474 278
TriviaQA multi-hop 78,785 8,837
HotpotQA multi-hop 90,447 7,405
Table 1. Statistics of different datasets
We evaluate two adversarial loss strategies:
  • RAAT: Reduces performance gap between best and worst retrieval cases, ensuring stability under challenging conditions.
  • IRM: Minimizes performance variance across retrieval environments, mitigating sensitivity to distribution shifts for reliable cross-scenario performance.
3.3. Experimental Setups
Model RAG Methods HotpotQA NQ WebQuestions TriviaQA AVERAGE
@@ 3.3. Experimental Setups
Llama-2-7b-chat-hf
Base Model 3.30 12.34 1.21 10.61 0.00 13.08 4.32 20.27 2.21 14.08
@@
$\Delta$ (Worst $\rightarrow$ Best) 31.17% 29.35% 33.39% 24.63% 59.60% 34.39% 14.65% 17.66% 31.16% 24.72%
Llama-3-8B-Instruct
Base Model 23.31 32.60 30.04 41.59 26.98 43.25 58.80 66.45 34.78 45.97
RALM 30.73 41.25 35.19 46.10 47.84 56.98 54.75 63.27 41.36 51.11
RetRobust 36.88 49.43 43.28 55.04 52.88 62.10 59.06 67.77 48.03 58.59
Top-1 Doc 37.09 49.67 44.38 56.20 54.68 62.26 60.80 68.31 49.24 59.11
Golden Doc 35.68 48.97 41.35 53.13 48.92 58.41 58.26 66.99 46.05 56.88
Random Doc 36.73 49.53 43.37 55.43 53.24 62.55 60.62 68.64 48.49 59.04
Irrelevant Doc 35.53 48.27 42.45 54.41 46.76 57.67 58.97 66.57 45.93 56.73
RAAT 32.79 44.26 42.34 53.31 48.28 58.17 54.41 62.45 44.46 54.55
IRM 35.63 48.72 41.13 53.14 53.96 61.64 57.15 69.13 46.97 58.16
$\Delta$ (Worst $\rightarrow$ Best) 20.70% 20.41% 26.12% 21.91% 16.94% 9.78% 11.74% 10.70% 16.88% 13.89%
@@
Table 2. Performance comparison of Llama-2-7b-chat-hf and Llama-3-8B-Instruct across various robust RAG methods on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA). The row $\Delta$ (Worst $\rightarrow$ Best) indicates the marginal robustness benefit achieved by the best method compared to the worst method.
Model RAG Methods HotpotQA NQ WebQuestions TriviaQA AVERAGE
EM F1 EM F1 EM F1 EM F1 EM F1
Qwen1.5-7B-Chat
Base Model 24.15 34.06 24.31 35.27 23.02 38.65 48.43 57.65 29.98 41.41
RALM 24.44 34.23 29.83 40.22 44.24 53.46 46.40 55.96 36.23 45.97
RetRobust 29.05 40.99 34.82 45.92 47.12 55.39 49.16 59.01 40.04 50.33
Top-1 Doc 29.68 41.57 34.16 45.44 43.88 52.71 49.61 59.03 39.33 49.69
Golden Doc 28.66 40.55 32.93 43.99 43.88 52.74 49.25 58.87 38.68 49.04
Random Doc 28.83 40.66 33.96 45.09 46.04 53.21 48.75 58.54 39.40 49.38
IRM 26.95 38.67 30.59 41.83 47.84 55.68 45.52 56.24 37.73 48.11
Qwen2.5-7B-Instruct
Base Model 25.10 35.02 26.97 38.15 25.90 42.46 53.86 62.73 32.96 44.59
RALM 26.37 36.30 31.43 42.00 41.73 51.98 53.92 62.74 38.36 48.26
RetRobust 30.47 42.26 34.60 46.00 45.68 53.80 54.43 63.63 41.30 51.42
Top-1 Doc 30.24 42.61 34.72 46.20 46.04 54.13 53.65 63.01 41.16 51.49
Golden Doc 30.28 41.84 33.77 45.11 44.60 53.18 54.35 63.55 40.75 50.92
Random Doc 30.25 42.22 34.08 45.47 44.24 53.30 54.29 63.54 40.72 51.13
IRM 29.58 41.73 32.79 44.17 46.40 54.80 53.45 64.26 40.56 51.24
Table 3. Performance comparison of different LLMs (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) across different robust RAG scenarios on four datasets (HotpotQA, NQ, WebQuestions, and TriviaQA).
Datasets
For our experiment, we evaluate on four widely-used question answering datasets: (1) single-hop QA, including NaturalQuestions (NQ) (Kwiatkowski et al., 2019) and WebQuestions (Berant et al., 2013); and (2) multi-hop QA, including TriviaQA (Joshi et al., 2017) and HotpotQA (Yang et al., 2018). All experimental results are evaluated on their dev splits using the Exact Match (EM) and F1 metrics. Detailed statistics of these datasets are listed in Table 1.
Implementation Details
We implement a standard RAG framework with retrieval and generation phases. For retrieval, we use Contriever (Izacard et al., 2022), a BERT-based dense retriever to retrieve top-20 documents per query as a candidate pool. We sample different document subsets from the candidate pool to create varying retrieval environments for each methods. For question-answering, the top-5 retrieved documents are concatenated with the query as input for answer generation. For model training, we utilize the LLaMA-Factory library to facilitate efficient LLM finetuning. We train our models with a learning rate of 1e-6 and set the maximum number of epochs to 3. To optimize the training process, we employ bfloat16 precision and DeepSpeed ZeRO-3 for distributed training. The gradient accumulation steps are set to 2 with a batch size of 8. For inference, we leverage vLLM to ensure efficient model serving. We maintain the default decoding parameters for each model during inference, including the default temperature and top-p sampling probabilities, to ensure fair comparison across different model variants. All training and inference procedures are conducted on 8 $\times$ NVIDIA A100 80GB GPUs.
4. Diminishing Benefits of Complex Methods With Model Capacity